With a tail wind, a plain flew 1740 km in 6 hours. With a head wind, it flew only 570 km in 3 hours. Find the plains air speed and the wind speed. Assume both are constant.
Answer by checkley77(12844)

D=RT
1740=R*6
R=1740/6
R=290 KMH. WITH THE TAIL WIND.
570=R*3
R=570/3
R=190 KMH. AGAINST THE WIND.
(290-190)/2=100/2=50 KMH. IS THE HEAD WIND.
PROOF:
290-50=240 KMH. FOR THE RATE OF THE PLANE
190+50=240 KMH. DITTO.